### v2

# ShadowMaskFormer: Mask Augmented Patch Embedding for Shadow Removal

Abstract
Transformer recently emerged as the de facto model for computer vision tasks and has also been successfully applied to shadow removal. However, these existing methods heavily rely on intricate modifications to the attention mechanisms within the transformer blocks while using a generic patch embedding. As a result, it often leads to complex architectural designs requiring additional computation resources. In this work, we aim to explore the efficacy of incorporating shadow information within the early processing stage. Accordingly, we propose a transformer-based framework with a novel patch embedding that is tailored for shadow removal, dubbed ShadowMaskFormer. Specifically, we present a simple and effective mask-augmented patch embedding to integrate shadow information and promote the model’s emphasis on acquiring knowledge for shadow regions. Extensive experiments conducted on the ISTD, ISTD+, and SRD benchmark datasets demonstrate the efficacy of our method against state-of-the-art approaches while using fewer model parameters.
Our research introduces ShadowMaskFormer, a transformer-based framework designed to enhance shadow removal in images. This new approach simplifies the process and improves efficiency, requiring fewer model parameters compared to existing methods. Technologically, ShadowMaskFormer integrates shadow information early in the processing stage, enabling more accurate and less resource-intensive image analysis. This can lead to more cost-effective AI applications where computational resources or power efficiency is a concern. Economically, the reduction in computational demand may lower the barriers to implementing advanced image-processing technologies in consumer electronics and other devices. Socially, by improving the quality of shadow removal, our method could enhance the visual experience in applications such as digital photography and video, making these technologies more accessible and enjoyable for users. ShadowMaskFormer contributes to the ongoing development of AI in visual computing by offering a more streamlined approach to a common problem, potentially influencing future advancements in the field.
Patch Embedding, Shadow Mask, Shadow Removal, Vision Transformer, Deep Learning.
1 Introduction
Deep learning-based approaches have been widely used for various computer vision tasks by performing excellent performance over traditional model-based approaches [1, 2, 3, 4]. Shadow removal is highly challenging as it involves restoring irregular shadow regions and has gradually emerged as the dominant paradigm in this field. Recently, deep learning-based approaches [5, 6, 7, 8, 9] have gradually emerged as the dominant paradigm in this field. It should be emphasized that shadow mask which can distinguish shadow regions from non-shadow regions in an image, has been widely employed in deep learning methods and proven to effectively assist in shadow removal task [8, 10, 11].
Limitation of Shadow Transformer. Recently, as an emerging backbone model of choice for vision tasks, transformers have also been applied to shadow removal tasks. However, there are two issues. ❶ The existing transformer-based methods overlook the shadow information in the early processing stage and directly employ the generic patch embedding, which cannot fully unleash the representational power of the transformer-based methods. As shown in Figure 1(a), the state-of-the-art transformer-based methods (i.e., CRFormer [12] and ShadowFormer [13]) achieve the contextual knowledge from the non-shadow regions, resulting in performance degradation. ❷ As shown in Figure 1(c)(c), the transformer-based methods mainly necessitate the creation of new modules with shadow masks within the main computation blocks (i.e., transformer blocks and convolution blocks) for shadow removal, resulting in a noticeable scale of parameters. Hence, this has led to the following research question: Can we incorporate shadow information in patch embedding to avoid complicated modifications to transformer blocks and highlight the shadow region in the feature extraction stage?
ShadowMaskFormer. To overcome the constraints of current transformer-based techniques and address the research question, we introduce a new framework called ShadowMaskFormer. As shown in Figure 1(b), ShadowMaskFormer combines the transformer model with a shadow mask in patch embedding, presenting a novel approach to effectively remove shadows from images. Specifically, in the early processing stage (i.e., patch embedding), we propose a simple and effective patch embedding, namely Mask Augmented Patch Embedding (MAPE). The motivation behind MAPE is our observations based on the limited utilization of masks from existing work, as detailed in Section 3.2. Therefore, in our ShadowMaskFormer, the shadow mask is carefully utilized in MAPE with two complementary binarization schemes (the and Binarization) to enhance the shadow region pixels, as detailed in section 4.2. In conclusion, with the proposed MAPE, our ShadowMaskFormer can not only leverage the contextual information that is achieved by the learning capability of the transformer model but also restore the shadow region pixels at an earlier time. This approach allows the model to more effectively acquire the distinctive features of shadow regions and perform shadow removal with more purposefully. It is noteworthy that MAPE operates as a single-layer module, primarily performing pixel-level operations, and it requires computation only once per training epoch, significantly reducing computational demands. As far as our knowledge extends, we believe our approach marks the pioneering exploration of utilizing patch embedding in vision transformer models for the task of shadow removal. Experimental results demonstrate that ShadowMaskFormer achieves outstanding shadow removal results over the three widely-used shadow removal datasets, surpassing the state-of-the-art performance. Furthermore, compared with other SOTA methods, it only has 2.2MB network parameters, e.g., half of CRFormer [12] and smaller than others, as illustrated in Figure 1(c). The main contributions of this work are as follows:
We propose an innovative transformer-based framework called ShadowMaskFormer111Our implementation is available at https://www.github.com/lizhh268/ShadowMaskFormer.git, which incorporates mask information in the patch embedding stage. To the best of our knowledge, we are the first to explore the task of shadow removal from the perspective of patch embedding in vision transformer models.
We propose the concept patch embedding tailored for shadow removal and further introduce the corresponding MAPE which is simple and effective. By carefully utilizing the shadow mask with two complementary binarization schemes, MAPE effectively enhances the shadow region to assist in restoring each pixel without introducing any model parameters.
Comprehensive experiments conducted on publicly available datasets, namely ISTD, ISTD+, and SRD, demonstrate the outstanding performance of our proposed method, placing it at the state-of-the-art level.
The remainder of this paper is structured as follows: Section II reviews related work in shadow removal and vision transformer. Section III introduces the task background and the motivation of our proposed method. Section IV describes the detailed architecture and workflow of ShadowMaskFormer. Sections V and VI present our method’s experimental results and ablation studies. Finally, Section VII concludes our work and discusses future directions.
2 Related Work
In this section, We will present prior works related to image shadow removal, briefly outline vision transformers, and discuss their application in the task of shadow removal.
2.1 Shadow Removal
In the field of image processing, shadows pose a pervasive challenge that has negative implications for various downstream tasks, e.g., object detection, tracking, and face recognition [14, 15, 16]. Consequently, shadow removal has been a fundamental task in the field of computer vision and has received extensive research attention over the years. Numerous methods have emerged for shadow removal in images. These methods can be broadly categorized into two flavors: traditional model-based approaches and deep learning-based approaches. Traditional model-based approaches rely on the physical models of shadow images, which are limited by their dependence on prior knowledge and often struggle to effectively remove shadows in real-world scenes [17, 18, 19]. For instance, [17] proposed a physical model based on constant illumination conditions; [20] proposed an optimization model for removing shadow under various illumination conditions.
In the recent past, deep learning-based (DL) methods have achieved excellent performance in the field of shadow removal, leveraging their end-to-end capabilities.
DL methods can be primarily categorized into two main approaches: convolutional neural network (CNN) based and generative modeling, particularly generative adversarial network (GAN), based approaches. CNN-based approaches employ multi-level convolutions to extract context-based features for shadow removal [5, 7, 21, 22, 23, 24, 6, 25, 26]. For instance, DeshadowNet [5] utilizes image contextual information to learn shadow mask features and remove shadows. GAN-based approaches generate shadow mask images and shadow-free images that closely resemble reality by adhering to a series of criteria for discerning between real and fake data [9, 11, 8, 10, 27, 28, 29, 28]. For instance, in DC-ShadowNet [28], three dedicated losses were defined to characterize shadow images for accurate shadow synthesis. Besides, the most recent work [30] predicts the adaptive weights between two features that are extracted from the shadow image and the shadow mask image. Our method is different from them, i.e., we define the problem as the integration of transformer models and shadow masks, which enables us to draw inspiration from existing mask utilization methods.
2.2 Vision Transformer
The transformer model was originally designed for natural language processing tasks [31], which were recently adopted for vision tasks. Specifically, vision transformers segment images into patches and utilize them as inputs for the subsequent computation (i.e., multi-head attention) [32]. Owing to the ability to learn global contextual information, transformers have achieved remarkable performance in vision tasks [33, 34]. Apart from regular image classification [33] and segmentation [35], transformer models have also been applied to various low-level vision tasks, such as image restoration [36], colorization [37], and inpainting [38]. Differing from other image tasks, our objective is to propose a transformer-based framework for shadow removal. This allows us to design a module for utilizing shadow masks, achieving efficient shadow removal.
2.3 Transformer for Shadow Removal
With the remarkable performance of transformer models in computer vision tasks, it has also been applied to shadow removal tasks [12, 13]. CRFormer [12] attempts to guide the restoration of shadow regions by leveraging non-shadow region information with shadow masks. However, it places excessive emphasis on non-shadow region information, resulting in insufficient model attention allocation to the crucial shadow regions and shadow masks. ShadowFormer [13] proposes a Shadow-Interaction Module Attention to exploit the global contextual correlation between shadow and non-shadow regions. It is worth noting that the transformer-based method, still like other DL-based methods, relies on designing new modules from the transformer blocks to leverage shadow masks for the shadow removal task, as shown in Figure 1(a). In this work, we aim to explore the efficacy of incorporating shadow information during the early processing stage and propose a novel patch embedding module tailored to the shadow removal task.
3 Preliminaries
In this section, we will provide the essential background knowledge about DL models and image shadow removal. Furthermore, we will elucidate the shadow model in this work and its motivation.
3.1 Background
In conventional computer vision models, such as Convolutional Neural Networks (CNNs), an input image is processed as an entirety by sliding a filter over the entire image to capture local patterns. In contrast, a vision transformer (ViT) model [32] does not directly operate on an input image, rather, it breaks down an input image into smaller and non-overlapping patches. Each patch is then linearly transformed into a fixed-dimensional vector representation for subsequent computation. This process is known as Patch Embedding.
The main computing backbone of a ViT model comprises sequentially connected transformer blocks. Each block comprises a multi-head attention (MHA), a feed-forward network (FFN), and LayerNorms [39].
3.2 Motivation
Shadow masks. With a steady stream of promising empirical results confirming the effectiveness of shadow masks for shadow removal [8, 11, 12, 13], we observe two patterns emerged from existing works: ❶ Shadow masks are typically represented as binary masks (i.e., with being the size of an input image) where “” indicates shadow regions and “” indicates non-shadow regions, respectively [8, 23, 12, 13]; ❷ Shadow masks are primarily utilized by the main computing units of the models, e.g., convolution operations in CNNs [40, 23] and attention modules in ViTs [12, 13].
Regarding observation ❶, binarization to poses a potential risk of losing useful information as features corresponding to non-shadow regions will be completely suppressed if one directly applies M to input signals. Instead of designing alternative ways to indirectly apply M as done in most prior arts [8, 23, 12, 13], we opt for designing two complementary binarization schemes for direct utilization of M in this work.
In response to observation ❷, we re-visit the location within a deep learning model where M should be incorporated. Specifically, we seek to investigate the efficacy of incorporating M during the input preprocessing stage, i.e., the patch-embedding stage for ViT models, eliminating the need for repeated applications of M in every transformer block, which in turn leads to improved model efficiency.
Physical models of shadow removal. For the model of shadow removal, based on the preceding works [41, 42, 43], it can be deduced that the formation of shadows occurs due to obstruction of direct illumination and a portion of the ambient illumination. This implies that the shadowed pixel will exhibit diminished intensity compared to their corresponding shadow-free pixels . According to [21, 43], we start from the original shadow illumination model, which describes a mapping function that transforms a shadow pixel to its non-shadow pixel . This mapping can be summarized as a linear function and the intensity of a lit pixel is formulated as:
where is the wavelength and is the intensity reflected from an image pixel, and are the illumination and reflectance respectively, and denote the direct illumination and the ambient illumination, respectively. A more detailed implementation of the shadow physics model can be found in the Appendix. For a real shadow scene, an occluder blocks the direct illumination and part of the ambient illumination , thus the shadowed pixel can be represented as:
where is the attenuation factor indicating the remaining fraction of that arrives at an image point .
From these Equations, the shadow-free pixel can also be expressed as follows:
Furthermore, following [21], we can establish a mapping function between and . This mapping expresses the shadow-free pixel as a linear function of the shadowed pixel.
where represents the color channel ( ,,), is the response of the camera to direct illumination, and is responsible for the attenuation factor of the ambient illumination at this pixel in this color channel. Additionally, and are constant across all pixels in the umbra area of the shadow. The crucial aspect of the learned mapping by the model lies in determining the parameters and for individual shadowed pixels. Inspired by this, we further assume the expression of Eq. 4 can be reformulated as follows:
In the above expression, the gain factor in our method is derived from the assumption underlying Eq. 1, and will be learned by the model for the shadowed pixel . Specifically, considering the natural attribute of shadows where the attenuation of light sources results in shadow pixel values significantly lower than those of non-shadowed pixels, we assume that in Eq. 1, will be noticeably greater than 1, and will be smaller compared to . The term appears to be roughly equivalent to the body reflection. Hence, the crux of shadow removal lies in ensuring that the gain factor learned by the model closely approximates the correct solution. In this paper, we endeavor to perform a preliminary exploration of the gain factor during the patch embedding stage. This approach allows the model to enhance shadow removal performance more effectively at an early stage, thereby reducing unnecessary model exploration. Specifically, to address the task of shadow removal, we introduce a simple and effective patch embedding, namely Mask Augmented Patch Embedding (MAPE).
4 ShadowMaskFormer
In this section, we will elucidate the overarching workflow framework of the model, with a particular emphasis on detailing the implementation of MAPE.
4.1 Overview
An overview of the proposed transformer with mask augmented in the patch embedding stage (ShadowMaskFormer) is depicted in Figure 2. This approach employs two types of shadow masks to enhance the shadow region pixels. To achieve this, in the early processing stage of the model, ShadowMaskFormer integrates a Mask Augmented Patch Embedding (MAPE) specifically designed for shadow removal.
Specifically, in the patch embedding stage, MAPE takes the shadow image and its corresponding shadow mask M as inputs. The main idea behind MAPE is to enhance the shadow region pixels in the early stage by utilizing the information of the M and . In ShadowMaskFormer, the determination of is achieved through two forms of mask utilization: initialization at the early stage and adaptive refinement through model exploration. After MAPE, the transformer blocks, based on the vision transformer, leverage its powerful ability to learn contextual information. It aims to learn a nonlinear mapping function from to . By applying this mapping function, the model reconstructs the shadow-free image J which represents the image after shadow removal.
4.2 Mask Augmented Patch Embedding
Based on the discussions above, aiming to find an accurate approximation of the gain factor for shadow region pixels early in the model’s training, we introduce the concept of Mask Augment Patch Embedding (MAPE) which is based on the two complementary binarization schemes (the and Binarization) as depicted in Figure 2. In particular, to ensure the learned non-linear mapping function of the model to explore the correct early and avoid exploring useless information for shadow region enhancement, we designed MAPE to be positioned in the early stage of the model. In the patch embedding stage, the shadow image and its corresponding shadow mask M are first used as inputs to the model. Subsequently, the and (Binarize) pixel-wise operations are performed on M, resulting in two complementary masks and , respectively. These operations can be expressed using the following formulas:
After applying binarization, pixels corresponding to the shadow regions of are set to 1, while the non-shadow regions are set to 0. undergoes pixel-wise operations to adjust the pixel distribution to the range [-1, 1].
First, we use a set of masks to learn weighted masks for adjusting the intensity of pixels from the shadow (S) and non-shadow (NS) regions, which in essence provides an initial estimation of . Thus we can obtain the first enhanced shadow image . Subsequently, the region information carried by is combined with through further operations, yielding the feature with element-wise multiplication. Finally, we employ linear projection on the , introducing linear transformation features. These operations of mask-augmented patch embedding can be expressed as follows.
where and are the weights for region reassignment. F is the output of MAPE, and denotes the convolutional layer with the kernel size of 3 × 3. Note that due to performance and efficiency by introducing convolutions into ViT [44], we opted for convolutional projection rather than positional projection.
Actually, the process of mask augmented can be seen as two steps and its pseudocode can be summarized as Algorithm 1. In the first step, we reassign the pixel information in the shadow and non-shadow regions by using the region reassignment weights and with mask, respectively. This can be expressed as , which means . It should be noted that in order to enhance the shadow region pixels, it is necessary to set greater than . In our experiments, we empirically set and to 2.5 and 1, respectively. Besides, we conducted a feature analysis of shadow images, e.g.: For ISTD datasets, the results indicate that after the processing, the proportion of positive values in the non-shadow regions reaches 70%. In contrast, in the shadow regions, negative values overwhelmingly dominate, reaching 97% of the total values. As shadow region pixel values tend to be low, the majority of shadow region pixels of are negative and this contradicts with the pixels in the non-shadow region (positive). Therefore, the second step is introduced to further enhance shadow region pixels by adaptive refinement. We use masks to re-balance the distribution of the normalized pixel intensity from the shadow and non-shadow regions, where most shadow pixels are negative-valued while most non-shadow pixels are positive-valued. This step essentially aims to make the restored shadow regions closely resemble non-shadow regions. Finally, we applied linear projection to , transforming it into the input for the transformer block. At this point, . Note that the transformer blocks also simultaneously explore the genuine values for each pixel during the training phase. In addition, MAPE avoids intricate modifications to the transformer blocks and eliminates the involvement of shadow masks in the computation of every transformer block, leading to noticeable computational savings.
4.3 Loss Function
5 Experiments
This section briefly describes our experimental setup followed by results on multiple shadow removal benchmarks.
| Method | Params | S | NS | All | |
| (M) | RMSE | RMSE | RMSE | ||
| 256 256 | Input Image | - | 32.1 | 7.09 | 10.9 |
| Guo et al. [45] | - | 18.7 | 7.76 | 9.26 | |
| MaskShadow-GAN [11] | 13.8 | 12.7 | 6.68 | 7.41 | |
| ST-CGAN [8] | 31.8 | 9.99 | 6.05 | 6.65 | |
| DSC [7] | 22.3 | 8.72 | 5.04 | 5.59 | |
| G2R [29] | 22.8 | 10.7 | 7.55 | 7.85 | |
| DHAN [10] | 21.8 | 7.49 | 5.30 | 5.66 | |
| Fu et al. [6] | 187 | 7.91 | 5.51 | 5.88 | |
| DC-ShadowNet [28] | 21.2 | 11.4 | 5.81 | 6.57 | |
| Zhu et al. [23] | 10.1 | 8.29 | 4.55 | 5.09 | |
| CRFormer† [12] | 4.89 | 7.32 | 5.82 | 6.07 | |
| ShadowFormer [13] | 2.40 | 6.16 | 3.90 | 4.27 | |
| MAPE (ours) | 2.28 | 6.08 | 3.86 | 4.23 | |
| 480 640 | Input Image | - | 33.23 | 7.25 | 11.4 |
| ARGAN [9] | - | 9.21 | 6.27 | 6.63 | |
| DHAN [10] | 21.8 | 8.13 | 5.94 | 6.29 | |
| CANet [22] | 358 | 8.86 | 6.07 | 6.15 | |
| ShadowFormer [13] | 2.40 | 6.93 | 4.59 | 4.96 | |
| MAPE (ours) | 2.28 | 6.83 | 4.50 | 4.87 |
| Method | Params | S | NS | All |
| (M) | RMSE | RMSE | RMSE | |
| Input Image | - | 40.2 | 2.6 | 8.5 |
| DHAN [10] | 21.8 | 11.2 | 7.1 | 7.8 |
| Param-Net [25] | - | 9.7 | 2.9 | 4.1 |
| G2R [29] | 22.8 | 8.8 | 2.9 | 3.9 |
| SP+M-Net [21] | 141 | 7.9 | 2.8 | 3.6 |
| Fu et al. [6] | 187 | 6.6 | 3.8 | 4.2 |
| SG-ShadowNet† [24] | 6.20 | 5.9 | 2.9 | 3.4 |
| ShadowFormer [13] | 2.40 | 5.4 | 2.4 | 2.8 |
| AdaptiveFusionNetwork [30] | 23.9 | 5.9 | 2.9 | 3.4 |
| MAPE (ours) | 2.28 | 5.4 | 2.2 | 2.7 |
5.1 Experimental Setup
Datasets. We evaluate our method on three widely-used benchmarks for shadow removal: (1) ISTD dataset [8] comprises 1,870 image triplets (i.e., shadow images, shadow-free images, and shadow masks), from which 1,330 and 540 triplets are used for training and testing, respectively; (2) Adjusted ISTD (ISTD+) dataset [21] reduces illumination inconsistency between shadow and shadow-free images using an image processing algorithm while retaining the same number of triplets as the original ISTD dataset; (3) SRD dataset [5] comprises 2,680 training and 408 testing pairs of shadow and shadow-free images w/o ground truth shadow masks. Accordingly, we adopt the predicted masks provided by [10] for training and evaluation.
Baselines. Towards obtaining a representative evaluation of our method, we consider a wide range of baselines, including a traditional method [45], generative modeling-assisted approaches [11, 8, 29, 10, 9, 30], two transformer-based approaches [12, 13], among others [7, 28, 22].
Evaluation metrics. We use root mean square error (RMSE) between the estimated shadow-free images and the ground truth in the LAB color space, where lower values indicate better results. We also use the peak signal-to-noise ratio (PSNR) and the structural similarity (SSIM) [46] metrics to quantitatively compare the performance of various methods in the RGB color space, where higher values indicate better results.
Implementation details. We implement our method on top of a recent image restoration framework [47] with swin transformer [48]. For training, we use AdamW [49] optimizer with a batch size of one and an initial learning rate of , which is annealed to zero in 300 epochs following the cosine schedule [50]. All experiments are carried out on two NVIDIA GeForce GTX 3090 GPUs.
Method Params Shadow Region (S) Non-Shadow Region (NS) All Image (All) (M) PSNR SSIM RMSE PSNR SSIM RMSE PSNR SSIM RMSE Input Image - 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05 DeshadowNet [5] - - - 11.78 - - 4.84 - - 6.64 DSC [7] 22.3 30.65 0.960 8.62 31.94 0.956 4.41 27.76 0.903 5.71 DHAN [10] 21.8 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67 Fu et al. [6] 187 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50 DC-ShadowNet [28] 21.2 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65 CANet [22] 358 - - 7.82 - - 5.88 - - 5.98 SG-shadowNet [24] 6.20 - - 7.53 - - 2.97 - - 4.23 Zhu et al. [23] 10.1 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79 CRFormer [12] 4.89 - - 7.14 - - 3.15 - - 4.25 ShadowFormer [13] 2.40 36.13 0.988 6.05 35.95 0.986 3.55 32.38 0.955 4.09 MAPE (ours) 2.28 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
5.2 Experimental Results
Results on ISTD. Table 1 summarizes the results on the ISTD dataset. In general, we observe that our method consistently outperforms all considered baselines on different regions (i.e., shadow and non-shadow regions) under both input resolution settings (i.e., and ) with fewer model parameters. In particular, our method achieves 1.6 lower RMSE (averaged over whole images) with 9.3 fewer parameters than DC-ShadowNet [28]. Our method also achieves 1.4 lower RMSE (averaged over whole images) with 2.1 fewer parameters than CRFormer [12]. Our method is also all-around competitive against ShadowFormer [13].
Additional qualitative comparisons are provided in Figure 3. In general, the empirical improvements provided by our method (as depicted in Table 1) translate well to visualization results. Evidently, we observe that our method significantly outperforms existing approaches in removing shadows from complex scenes (i.e., the second row), achieving a more natural adaptation between shadow and non-shadow regions. For example, SP+M-Net [21] tends to incorrectly process the non-shadow regions excessively, e.g., the results of the second and third rows in the fourth column. DC-ShadowNet [28] fails to successfully remove shadows and also affects the non-shadow regions, as shown in the result of the third row in the third column.
Results on ISTD+. Table 2 summarizes the results on the ISTD+ dataset. In general, we observe that (i) all methods exhibit better performance than that on the original ISTD dataset [8] after adjustments of illumination inconsistency introduced by [21]; (ii) our method continues to outperform all the considered baseline competitors as indicated by RMSE while using fewer model parameters. Specifically, our method achieves better RMSE on shadow, non-shadow, and all regions than both SG-ShadowNet [24] and ShadowFormer [13].
Results on SRD. Table 3 presents the comparison results on the SRD [5] dataset. Evidently, we observe that our method achieves significantly better results on the overall image than all considered baselines. For instance, our method yields 1.58dB to 2.28dB PSNR improvements on shadow, non-shadow, and all regions over ShadowFormer [13] with a similar number of model parameters. In addition, we observe that our method also achieves noticeably better RMSE than the baselines. Specifically, our method achieves 1.3 and 1.2 lower RMSE than CRFormer [12] on shadow and all regions, respectively, while using 2.1 fewer model parameters.
Figure 4 depicts qualitative comparisons between our method and baselines. Evidently, we observe that our method not only can better eliminate artifacts while removing shadows but also effectively restore the content and color of the shadow regions.
6 Ablation Studies
In this section, we conduct a detailed analysis of each component of our MAPE framework, with the number of transformer blocks N set to 5. We conduct ablation studies on the ISTD dataset [8] to evaluate the individual impact of these components. In our paper, we argue that the mask deep fusion approach efficiently utilizes shadow mask and provides useful and sufficient information about shadow regions.
Ablation Study on MAPE. As summarized in Table 4, we first present the performance of the regular path embedding [51], which is the de-facto choice of image preprocessing adopted by most existing works that are based on transformers. Evidently, our mask augmented patch embedding (MAPE) leads to significantly better shadow removal performance measured by RMSE. Compared to the conventional patch embedding, our MAPE can obviously utilize the regional information provided by the shadow mask effectively and convey this information to the entire model. Then, we present the relative improvements from the proposed binarization scheme by replacing it with the regular 0/1 binary mask (i.e., ), where the effectiveness of can be observed. This experimental result demonstrates that the combination of the two shadow masks designed in our MAPE is significantly more effective in shadow removal compared to using only a single, common 0/1 mask (). In other words, the success of MAPE relies on the reliable combined use of the two masks. On the contrary, applying shadow masks directly within patch embedding can lead to performance degradation in shadow regions, as shown in Table 4.
Method S NS All RMSE RMSE RMSE Original PE 7.59 (+1.5) 4.91 (+1.1) 5.35 (+1.1) MAPE 6.08 (+0.0) 3.86 (+0.0) 4.23 (+0.0) in MAPE 8.03 (+2.0) 4.49 (+0.6) 4.95 (+0.7)
The effect of the Model’s variants. Our method, MAPE, is built upon the vision-transformer model and is sensitive to different model configurations. To investigate the impact of varying transformer block settings, we conducted experiments with three distinct model configurations on the ISTD dataset. The results are summarized in Table 5. We observed that increasing the number of model parameters enhances shadow removal capabilities, leading to a significant improvement from 1.94MB to 2.52MB in overall image performance. While higher model capacity improves performance, it also increases computational time, which can be burdensome. To address this concern, we selected the middle model parameter configuration to strike a balance between performance and computational efficiency, alleviating the overhead.
Model Size Params S NS All (M) RMSE RMSE RMSE Small 1.94 6.61 3.93 4.37 Middle 2.28 6.08 3.86 4.23 Large 2.52 6.27 3.80 4.20
7 Conclusion
In this paper, we introduce ShadowMaskFormer, a novel early mask utilization framework based on the transformer model, for efficient shadow removal. In ShadowMaskFormer, we successfully explore a simple and effective implementation of enhanced shadow region pixels during the patch embedding stage, enabling the model to restore shadow images. Specifically, after analyzing the characteristics of the previous methods, we develop the Mask Augmented patch embedding by using carefully the shadow mask. This approach efficiently encourages the model to explore the optimal shadow pixel gain factors as soon as possible. Experimental results on the ISTD, ISTD+, and SRD datasets demonstrate that our ShadowMaskFormer achieves outstanding performance compared to state-of-the-art methods by using fewer network parameters. In the future, we will further explore more shadow removal methods based on the physical parameters of the shadow model, to enhance the interpretability of the model and its adaptability to different scenes.

### v3

@@
 Extensive experiments conducted on the ISTD, ISTD+, and SRD benchmark datasets demonstrate the efficacy of our method against state-of-the-art approaches while using fewer model parameters.
+Our implementation is available at https://github.com/lizhh268/ShadowMaskFormer.
 Our research introduces ShadowMaskFormer, a transformer-based framework designed to enhance shadow removal in images.
@@
 As shown in Figure 1(a), the state-of-the-art transformer-based methods (i.e., CRFormer [12] and ShadowFormer [13]) achieve the contextual knowledge from the non-shadow regions, resulting in performance degradation.
-❷ As shown in Figure 1(c)(c), the transformer-based methods mainly necessitate the creation of new modules with shadow masks within the main computation blocks (i.e., transformer blocks and convolution blocks) for shadow removal, resulting in a noticeable scale of parameters.
+❷ As shown in Figure 1(c), the transformer-based methods mainly necessitate the creation of new modules with shadow masks within the main computation blocks (i.e., transformer blocks and convolution blocks) for shadow removal, resulting in a noticeable scale of parameters.
 Hence, this has led to the following research question: Can we incorporate shadow information in patch embedding to avoid complicated modifications to transformer blocks and highlight the shadow region in the feature extraction stage?
@@
 In conclusion, with the proposed MAPE, our ShadowMaskFormer can not only leverage the contextual information that is achieved by the learning capability of the transformer model but also restore the shadow region pixels at an earlier time.
-This approach allows the model to more effectively acquire the distinctive features of shadow regions and perform shadow removal with more purposefully.
+This approach allows the model to more effectively acquire the distinctive features of shadow regions and perform shadow removal more purposefully.
 It is noteworthy that MAPE operates as a single-layer module, primarily performing pixel-level operations, and it requires computation only once per training epoch, significantly reducing computational demands.
@@
 Experimental results demonstrate that ShadowMaskFormer achieves outstanding shadow removal results over the three widely-used shadow removal datasets, surpassing the state-of-the-art performance.
-Furthermore, compared with other SOTA methods, it only has 2.2MB network parameters, e.g., half of CRFormer [12] and smaller than others, as illustrated in Figure 1(c).
+Furthermore, compared with other SOTA methods, it only has 2.2MB network parameters, e.g., half of CRFormer [12] and smaller than others, as illustrated in Figure 1.
 The main contributions of this work are as follows:
-We propose an innovative transformer-based framework called ShadowMaskFormer111Our implementation is available at https://www.github.com/lizhh268/ShadowMaskFormer.git, which incorporates mask information in the patch embedding stage.
-To the best of our knowledge, we are the first to explore the task of shadow removal from the perspective of patch embedding in vision transformer models.
-We propose the concept patch embedding tailored for shadow removal and further introduce the corresponding MAPE which is simple and effective.
-By carefully utilizing the shadow mask with two complementary binarization schemes, MAPE effectively enhances the shadow region to assist in restoring each pixel without introducing any model parameters.
-Comprehensive experiments conducted on publicly available datasets, namely ISTD, ISTD+, and SRD, demonstrate the outstanding performance of our proposed method, placing it at the state-of-the-art level.
+We propose ShadowMaskFormer, a novel framework that integrates shadow mask information in the patch embedding stage.
+To the best of our knowledge, this is the first attempt to incorporate shadow-specific information at such an early stage in transformer models, establishing a new paradigm for shadow removal tasks.
+We present MAPE, a simple yet effective method tailored for integrating shadow information, naturally aligned with the patch embedding stage.
+By leveraging shadow masks with two complementary binarization schemes, MAPE enhances shadow region representations, enabling precise pixel-wise restoration without adding extra parameters.
+This results in both computational efficiency and robust performance across a wide range of scenarios.
+Our method achieves state-of-the-art performance across benchmark datasets (ISTD, ISTD+, and SRD), demonstrating both exceptional accuracy and computational efficiency with a model size of only 2.2MB.
+Extensive experiments further demonstrate the robustness of MAPE even under incomplete or inaccurate shadow masks, proving its outstanding ability to generalize to complex, real-world conditions.
 The remainder of this paper is structured as follows: Section II reviews related work in shadow removal and vision transformer.
@@
 For instance, [17] proposed a physical model based on constant illumination conditions; [20] proposed an optimization model for removing shadow under various illumination conditions.
-In the recent past, deep learning-based (DL) methods have achieved excellent performance in the field of shadow removal, leveraging their end-to-end capabilities.
-DL methods can be primarily categorized into two main approaches: convolutional neural network (CNN) based and generative modeling, particularly generative adversarial network (GAN), based approaches.
+In the recent past, deep learning-based methods have achieved excellent performance in the field of shadow removal, leveraging their end-to-end capabilities.
+Deep learning-based methods can be primarily categorized into two main approaches: convolutional neural network (CNN) based and generative modeling, particularly generative adversarial network (GAN), based approaches.
 CNN-based approaches employ multi-level convolutions to extract context-based features for shadow removal [5, 7, 21, 22, 23, 24, 6, 25, 26].
@@
 For instance, in DC-ShadowNet [28], three dedicated losses were defined to characterize shadow images for accurate shadow synthesis.
-Besides, the most recent work [30] predicts the adaptive weights between two features that are extracted from the shadow image and the shadow mask image.
+Besides, AdaptiveFusionNetwork [30] predicts the adaptive weights between two features that are extracted from the shadow image and the shadow mask image.
+On the other hand, two recent works have made promising progress in shadow removal: ShadowDiffusion [31] achieves notable performance by proposing a dynamic mask-aware diffusion model and incorporating degradation priors, while HomoFormer [32] offers another solution by using random shuffle and reverse shuffle operations to transform non-uniform shadows into a uniform distribution, enhancing the effectiveness of local self-attention for complex shadows.
 Our method is different from them, i.e., we define the problem as the integration of transformer models and shadow masks, which enables us to draw inspiration from existing mask utilization methods.
@@
 ShadowFormer [13] proposes a Shadow-Interaction Module Attention to exploit the global contextual correlation between shadow and non-shadow regions.
-It is worth noting that the transformer-based method, still like other DL-based methods, relies on designing new modules from the transformer blocks to leverage shadow masks for the shadow removal task, as shown in Figure 1(a).
+It is worth noting that the transformer-based method, still like other deep learning-based methods, relies on designing new modules from the transformer blocks to leverage shadow masks for the shadow removal task, as shown in Figure 1(a).
 In this work, we aim to explore the efficacy of incorporating shadow information during the early processing stage and propose a novel patch embedding module tailored to the shadow removal task.
@@
 3.1 Background
-In conventional computer vision models, such as Convolutional Neural Networks (CNNs), an input image is processed as an entirety by sliding a filter over the entire image to capture local patterns.
+In conventional computer vision models, such as convolutional neural networks, an input image is processed as an entirety by sliding a filter over the entire image to capture local patterns.
 In contrast, a vision transformer (ViT) model [34] does not directly operate on an input image, rather, it breaks down an input image into smaller and non-overlapping patches.
@@
 Shadow masks.
-With a steady stream of promising empirical results confirming the effectiveness of shadow masks for shadow removal [8, 11, 12, 13], we observe two patterns emerged from existing works: ❶ Shadow masks are typically represented as binary masks (i.e., with being the size of an input image) where “” indicates shadow regions and “” indicates non-shadow regions, respectively [8, 23, 12, 13]; ❷ Shadow masks are primarily utilized by the main computing units of the models, e.g., convolution operations in CNNs [40, 23] and attention modules in ViTs [12, 13].
+With a steady stream of promising empirical results confirming the effectiveness of shadow masks for shadow removal [8, 11, 12, 13], we observe two patterns emerged from existing works: ❶ Shadow masks are typically represented as binary masks (i.e., with being the size of an input image) where “” indicates shadow regions and “” indicates non-shadow regions, respectively [8, 23, 12, 13]; ❷ Shadow masks are primarily utilized by the main computing units of the models, e.g., convolution operations in CNN-based models [42, 23] and attention modules in ViT models [12, 13].
 Regarding observation ❶, binarization to poses a potential risk of losing useful information as features corresponding to non-shadow regions will be completely suppressed if one directly applies M to input signals.
@@
 This approach allows the model to enhance shadow removal performance more effectively at an early stage, thereby reducing unnecessary model exploration.
-Specifically, to address the task of shadow removal, we introduce a simple and effective patch embedding, namely Mask Augmented Patch Embedding (MAPE).
+Specifically, to address the task of shadow removal, we introduce a simple and effective patch embedding, namely Mask Augmented Patch Embedding.
 4 ShadowMaskFormer
-In this section, we will elucidate the overarching workflow framework of the model, with a particular emphasis on detailing the implementation of MAPE.
+In this section, we will elucidate the overarching workflow framework of the model, with a particular emphasis on detailing the implementation of Mask Augmented Patch Embedding.
 4.1 Overview
@@
 4.2 Mask Augmented Patch Embedding
-Based on the discussions above, aiming to find an accurate approximation of the gain factor for shadow region pixels early in the model’s training, we introduce the concept of Mask Augment Patch Embedding (MAPE) which is based on the two complementary binarization schemes (the and Binarization) as depicted in Figure 2.
+Based on the discussions above, aiming to find an accurate approximation of the gain factor for shadow region pixels early in the model’s training, we introduce the concept of Mask Augmented Patch Embedding which is based on the two complementary binarization schemes (the and Binarization) as depicted in Figure 2.
 In particular, to ensure the learned non-linear mapping function of the model to explore the correct early and avoid exploring useless information for shadow region enhancement, we designed MAPE to be positioned in the early stage of the model.
@@
 4.3 Loss Function
+Following prior shadow removal works [21, 24, 12], we only adopt the image consistency loss , which is mathematically defined as follows:
+where and are the predicted shadow-free image and the ground truth, respectively.
+The use of loss is motivated by its ability to effectively capture pixel-wise differences, which is crucial for accurately removing shadows while preserving fine details.
+This makes loss well-suited to meet the requirements of our proposed framework.
 5 Experiments
@@
 | Method | Params | S | NS | All | |
-| (M) | RMSE | RMSE | RMSE | ||
+| (M) | MAE | MAE | MAE | ||
 | 256 256 | Input Image | - | 32.1 | 7.09 | 10.9 |
@@
 | CRFormer† [12] | 4.89 | 7.32 | 5.82 | 6.07 | |
+| ShadowDiffusion [31] | 55.52 | 4.13 | 4.14 | 4.12 | |
 | ShadowFormer [13] | 2.40 | 6.16 | 3.90 | 4.27 | |
-| MAPE (ours) | 2.28 | 6.08 | 3.86 | 4.23 | |
+| ShadowMaskFormer (ours) | 2.28 | 6.08 | 3.86 | 4.23 | |
 | 480 640 | Input Image | - | 33.23 | 7.25 | 11.4 |
@@
 | ShadowFormer [13] | 2.40 | 6.93 | 4.59 | 4.96 | |
-| MAPE (ours) | 2.28 | 6.83 | 4.50 | 4.87 |
+| ShadowMaskFormer (ours) | 2.28 | 6.83 | 4.50 | 4.87 |
 | Method | Params | S | NS | All |
-| (M) | RMSE | RMSE | RMSE | |
+| (M) | MAE | MAE | MAE | |
 | Input Image | - | 40.2 | 2.6 | 8.5 |
@@
 | SG-ShadowNet† [24] | 6.20 | 5.9 | 2.9 | 3.4 |
+| ShadowDiffusion [31] | 55.52 | 4.9 | 2.3 | 2.7 |
 | ShadowFormer [13] | 2.40 | 5.4 | 2.4 | 2.8 |
 | AdaptiveFusionNetwork [30] | 23.9 | 5.9 | 2.9 | 3.4 |
-| MAPE (ours) | 2.28 | 5.4 | 2.2 | 2.7 |
+| ShadowMaskFormer (ours) | 2.28 | 5.4 | 2.2 | 2.7 |
 5.1 Experimental Setup
@@
 Baselines.
-Towards obtaining a representative evaluation of our method, we consider a wide range of baselines, including a traditional method [45], generative modeling-assisted approaches [11, 8, 29, 10, 9, 30], two transformer-based approaches [12, 13], among others [7, 28, 22].
+Towards obtaining a representative evaluation of our method, we consider a wide range of baselines, including a traditional method [47], generative modeling-assisted approaches [11, 8, 29, 10, 9, 30], two transformer-based approaches [12, 13], among others [7, 28, 22, 31].
 Evaluation metrics.
-We use root mean square error (RMSE) between the estimated shadow-free images and the ground truth in the LAB color space, where lower values indicate better results.
+Following the previous methods [21, 25, 32], we use mean absolute error (MAE) between the estimated shadow-free images and the ground truth in the LAB color space, where lower values indicate better results.
 We also use the peak signal-to-noise ratio (PSNR) and the structural similarity (SSIM) [48] metrics to quantitatively compare the performance of various methods in the RGB color space, where higher values indicate better results.
@@
 All experiments are carried out on two NVIDIA GeForce GTX 3090 GPUs.
-Method Params Shadow Region (S) Non-Shadow Region (NS) All Image (All) (M) PSNR SSIM RMSE PSNR SSIM RMSE PSNR SSIM RMSE Input Image - 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05 DeshadowNet [5] - - - 11.78 - - 4.84 - - 6.64 DSC [7] 22.3 30.65 0.960 8.62 31.94 0.956 4.41 27.76 0.903 5.71 DHAN [10] 21.8 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67 Fu et al.
-[6] 187 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50 DC-ShadowNet [28] 21.2 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65 CANet [22] 358 - - 7.82 - - 5.88 - - 5.98 SG-shadowNet [24] 6.20 - - 7.53 - - 2.97 - - 4.23 Zhu et al.
-[23] 10.1 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79 CRFormer [12] 4.89 - - 7.14 - - 3.15 - - 4.25 ShadowFormer [13] 2.40 36.13 0.988 6.05 35.95 0.986 3.55 32.38 0.955 4.09 MAPE (ours) 2.28 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
+Method Params Shadow Region (S) Non-Shadow Region (NS) All Image (All) (M) PSNR SSIM MAE PSNR SSIM MAE PSNR SSIM MAE Input Image - 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05 DeshadowNet [5] - - - 11.78 - - 4.84 - - 6.64 DSC [7] 22.3 30.65 0.960 8.62 31.94 0.956 4.41 27.76 0.903 5.71 DHAN [10] 21.8 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67 Fu et al.
+[6] 187 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50 DC-ShadowNet [28] 21.2 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65 CANet [22] 358 - - 7.82 - - 5.88 - - 5.98 SG-shadowNet [24] 6.20 - - 7.53 - - 2.99 - - 4.23 Zhu et al.
+[23] 10.1 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79 CRFormer [12] 4.89 - - 7.14 - - 3.15 - - 4.25 ShadowDiffusion [31] 55.52 38.72 0.987 4.98 37.78 0.985 3.44 34.73 0.970 3.63 ShadowFormer [13] 2.40 36.13 0.988 6.05 35.95 0.986 3.55 32.38 0.955 4.09 ShadowMaskFormer (ours) 2.28 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
 5.2 Experimental Results
@@
 In general, we observe that our method consistently outperforms all considered baselines on different regions (i.e., shadow and non-shadow regions) under both input resolution settings (i.e., and ) with fewer model parameters.
-In particular, our method achieves 1.6 lower RMSE (averaged over whole images) with 9.3 fewer parameters than DC-ShadowNet [28].
-Our method also achieves 1.4 lower RMSE (averaged over whole images) with 2.1 fewer parameters than CRFormer [12].
+In particular, our method achieves 1.6 lower MAE (averaged over whole images) with 9.3 fewer parameters than DC-ShadowNet [28].
+Our method also achieves 1.4 lower MAE (averaged over whole images) with 2.1 fewer parameters than CRFormer [12].
 Our method is also all-around competitive against ShadowFormer [13].
@@
 Table 2 summarizes the results on the ISTD+ dataset.
-In general, we observe that (i) all methods exhibit better performance than that on the original ISTD dataset [8] after adjustments of illumination inconsistency introduced by [21]; (ii) our method continues to outperform all the considered baseline competitors as indicated by RMSE while using fewer model parameters.
-Specifically, our method achieves better RMSE on shadow, non-shadow, and all regions than both SG-ShadowNet [24] and ShadowFormer [13].
+In general, we observe that (i) all methods exhibit better performance than that on the original ISTD dataset [8] after adjustments of illumination inconsistency introduced by [21]; (ii) our method continues to outperform all the considered baseline competitors as indicated by MAE while using fewer model parameters.
+Specifically, our method achieves better MAE on shadow, non-shadow, and all regions than both SG-ShadowNet [24] and ShadowFormer [13].
 Results on SRD.
@@
 For instance, our method yields 1.58dB to 2.28dB PSNR improvements on shadow, non-shadow, and all regions over ShadowFormer [13] with a similar number of model parameters.
-In addition, we observe that our method also achieves noticeably better RMSE than the baselines.
-Specifically, our method achieves 1.3 and 1.2 lower RMSE than CRFormer [12] on shadow and all regions, respectively, while using 2.1 fewer model parameters.
+In addition, we observe that our method also achieves noticeably better MAE than the baselines.
+Specifically, our method achieves 1.3 and 1.2 lower MAE than CRFormer [12] on shadow and all regions, respectively, while using 2.1 fewer model parameters.
 Figure 4 depicts qualitative comparisons between our method and baselines.
@@
 As summarized in Table 4, we first present the performance of the regular path embedding [53], which is the de-facto choice of image preprocessing adopted by most existing works that are based on transformers.
-Evidently, our mask augmented patch embedding (MAPE) leads to significantly better shadow removal performance measured by RMSE.
+Evidently, our proposed mask augmented patch embedding leads to significantly better shadow removal performance measured by MAE.
 Compared to the conventional patch embedding, our MAPE can obviously utilize the regional information provided by the shadow mask effectively and convey this information to the entire model.
@@
 On the contrary, applying shadow masks directly within patch embedding can lead to performance degradation in shadow regions, as shown in Table 4.
-Method S NS All RMSE RMSE RMSE Original PE 7.59 (+1.5) 4.91 (+1.1) 5.35 (+1.1) MAPE 6.08 (+0.0) 3.86 (+0.0) 4.23 (+0.0) in MAPE 8.03 (+2.0) 4.49 (+0.6) 4.95 (+0.7)
+Method S NS All MAE MAE MAE Original PE 7.59 (+1.5) 4.91 (+1.1) 5.35 (+1.1) MAPE 6.08 (+0.0) 3.86 (+0.0) 4.23 (+0.0) in MAPE 8.03 (+2.0) 4.49 (+0.6) 4.95 (+0.7)
 The effect of the Model’s variants.
@@
 To address this concern, we selected the middle model parameter configuration to strike a balance between performance and computational efficiency, alleviating the overhead.
-Model Size Params S NS All (M) RMSE RMSE RMSE Small 1.94 6.61 3.93 4.37 Middle 2.28 6.08 3.86 4.23 Large 2.52 6.27 3.80 4.20
+Model Size Params S NS All (M) MAE MAE MAE Small 1.94 6.61 3.93 4.37 Middle 2.28 6.08 3.86 4.23 Large 2.52 6.27 3.80 4.20
+The effect of the quality of masks.
+Evaluating the robustness of MAPE to varying-quality shadow masks is crucial.
+We conducted a series of experiments and analyses to evaluate the robustness of MAPE to varying-quality shadow masks.
+Results suggest that MAPE maintains strong performance, even with inaccurate or low-quality masks.
+Specifically, the SRD dataset [5] uses shadow masks predicted by DHAN [10], which are incomplete and inaccurate, as shown in Figure 5.
+Despite this, our method achieves efficient shadow removal, demonstrating MAPE’s robustness to inaccurate input masks.
+To further quantify this robustness, we degraded the shadow masks in the ISTD dataset [8] to generate low-quality masks.
+Specifically, since the accuracy of shadow masks is related to the correctness of shadow boundaries, we first extracted the shadow region contours from the dataset’s shadow masks.
+We then randomly misclassified contour shadow pixels as non-shadow pixels to generate low-quality masks.
+As shown in the first row of Figure 6, different low-quality masks are presented, where the differences between low-quality and ISTD dataset masks are visible.
+The quality of these masks was evaluated using the balance error rate (BER) [54], where lower BER indicates higher accuracy.
+The results presented in Table 6 indicate that shadow removal performance is closely related to mask quality.
+When the BER increases, reflecting lower mask quality, the shadow removal performance degrades.
+However, even when the BER increases sevenfold (from 0.61 to 4.39), the MAE only changes by 11%, demonstrating that MAPE maintains its robustness to varying-quality masks.
+Finally, visual comparisons of shadow masks with different quality and the corresponding results are provided in Figure 6.
+We can observe slight contour issues in the shadow regions when BER is 0.61 or 1.82, and more pronounced issues when BER reaches 4.39.
+Mask S NS All BER MAE MAE MAE 0.61 7.13 4.19 4.67 1.29 7.36 4.21 4.72 1.82 7.78 4.20 4.78 2.85 8.47 4.22 4.93 3.36 8.87 4.24 5.02 4.39 9.72 4.29 5.21
+Generalization Performance of ShadowMaskFormer.
+To assess the generalization capability of ShadowMaskFormer, we conducted experiments on previously unseen scenarios.
+Specifically, we selected images from the ISTD dataset where shadows are cast on walls, an extremely common real-world scenario, as the unseen test set, while the remaining images were used for training.
+We maintained the same experimental setup as described in Section 5.1, ensuring consistency in the data partitioning strategy and evaluation approach.
+Unlike artificial or synthetic test settings, wall shadows naturally occur in diverse environments with textured surfaces and varying reflectance properties, providing a representative benchmark for evaluating the generalizability of shadow removal models in practical applications.
+The results, presented in Table 7, show that ShadowMaskFormer refers to the model trained and tested on the ISTD dataset, while ShadowMaskFormer* refers to the generalized results.
+We observe that ShadowMaskFormer* shows a slight decrease in performance in the shadowed regions, while performance in non-shadow regions improves, with overall image performance remaining consistent.
+These findings suggest that our method effectively adapts to unseen shadow scenarios, handling novel shadow characteristics while maintaining robust global performance.
+Methods S NS All MAE MAE MAE ShadowMaskFormer 6.08 3.86 4.23 ShadowMaskFormer* 6.69 3.55 4.23
+Additionally, as shown in Figure 7, the visual results further demonstrate that our method successfully removes shadows even in previously unseen settings.
+This highlights the strong generalization capability of ShadowMaskFormer, enabling it to handle real-world shadow removal tasks effectively.
 7 Conclusion